# From Limited Labels to Open Domains:An Efficient Learning Method for Drone-view Geo-Localization

### v3

@@ Abstract
Traditional supervised drone-view geo-localization (DVGL) methods heavily depend on paired training data and encounter difficulties in learning cross-view correlations from unpaired data. Moreover, when deployed in a new domain, these methods require obtaining the new paired data and subsequent retraining for model adaptation, which significantly increases computational overhead. Existing unsupervised methods have enabled to generate pseudo-labels based on cross-view similarity to infer the pairing relationships. However, geographical similarity and spatial continuity often cause visually analogous features at different geographical locations. The feature confusion compromises the reliability of pseudo-label generation, where incorrect pseudo-labels often drive negative optimization. Given these challenges inherent in both supervised and unsupervised DVGL methods, we propose a novel cross-domain invariant knowledge transfer network (CDIKTNet) with limited supervision, whose architecture consists of a cross-domain invariance sub-network (CDIS) and a cross-domain transfer sub-network (CDTS). This architecture facilitates a closed-loop framework for invariance feature learning and knowledge transfer. The CDIS is designed to learn cross-view structural and spatial invariance from a small amount of paired data that serves as prior knowledge. It endows the shared feature space of unpaired data with implicit cross-view correlations at initialization, which alleviates feature confusion. Based on this, the CDTS employs dual-path contrastive learning to further optimize each subspace while preserving consistency in a shared feature space. Extensive experiments demonstrate that CDIKTNet achieves state-of-the-art performance under full supervision compared with those supervised methods, and further surpasses existing unsupervised methods in both few-shot initialization and cross-domain initialization setting.
@@ I Introduction
Cross-view geo-localization has emerged as a promising visual localization technique [shore2024bev]. Early efforts primarily focus on retrieving ground-view images to geo-referenced satellite imagery for ground-view geo-localization [Cai_2019_ICCV]. With the widespread adoption of drones in complex scenarios such as emergency response [Liu_Multi] and urban surveys [Li_PVT++], the demand for autonomous localization in GPS-denied environments has been steadily increasing. This has driven more research interests in drone-view geo-localization (DVGL) [deuser2023sample4geo, wang2024multiple]. However, the dynamic viewpoints of drones result in scale distortions and non-rigid spatial transformations compared with satellite imagery. This pronounced structural and spatial heterogeneity presents challenges for DVGL. In recent years, benefiting from previous research efforts [wu2024camp, chen2024multi], supervised DVGL methods with paired data have made notable progress. These paired data often require manual labeling, which provides an accurate pairing relationship between satellite imagery and drone images of the same location.
@@
However, as illustrated in Fig. 1 (A(a)) and (B(a)), existing supervised methods [shen2023mccg, du2024ccr, xia2024enhancing] strictly rely on a large amount of paired data to learn cross-view correlations. The absence of sufficient paired data leads to a dramatic decline in model performance. When the supervised ratio drops to 0, or in other words, when there is no paired data and all the data are unpaired, the existing supervised methods become ineffective. We mark this situation with a red X in the figures. Furthermore, dynamic drone mission environments often lead to domain shifts, which in turn require repeated efforts to re-establish cross-view correlations before deployment in new domains. It hinders the applicability in scenarios where pairing relationships are difficult to obtain and cross-domain adaptation is required. As shown in Fig. 1 (B(a)), the new domain where all data are unpaired cannot directly employ the trained model from the source domain. Therefore, researchers have increasingly focused on unsupervised methods for DVGL. Recent work [11010141] constructed a sparse 3D point cloud from multi-view drone images and rendered pseudo-satellite images with near-orthographic views via 3D Gaussian splatting to realize geometric transformations of the drone images. These synthetic images are then matched against real satellite imagery to enable unsupervised training. However, this method is sensitive to 3D reconstruction and pose estimation errors, and it is difficult to bridge the inherent cross-view domain gaps. In contrast, another work [li2024learning] utilizes a frozen pre-trained backbone and a learnable adapter to align features from different views in shared feature space. The shared feature space serves as a unified latent representation in which drone and satellite views can be directly aligned. Cross-view pseudo-labels are generated automatically based on feature similarity as supervised signals to guide training. This method avoids explicit geometric transformations and aims to intrinsically learn cross-view correlations. Nevertheless, as shown in Fig. 2 (a), when visually similar patterns appear at different geographical locations, the resulting feature confusion in the shared feature space often leads to incorrect pseudo-label generation. This ultimately drives the model toward negative optimization.
@@
In this paper, we propose an efficient learning method with limited supervision for DVGL that mitigates the adverse effects of inaccurate pseudo-labels caused by feature confusion, called the cross-domain invariant knowledge transfer network (CDIKTNet). Specifically, the CDIKTNet consists of two core components, including the cross-domain invariance sub-network (CDIS) and the cross-domain transfer sub-network (CDTS). The CDIS utilizes a small amount of paired data to learn latent cross-view structural and spatial invariance, which serves as prior knowledge. As illustrated in Fig. 2(b), it endows the shared feature space of unpaired data with similar implicit cross-view correlations at initialization to alleviate feature confusion. Based on this, CDTS introduces dual-path contrastive learning to independently optimize subspaces under different views. Despite being optimized separately, these subspaces maintain alignment in the shared feature space due to the shared invariance prior. The main contributions can be summarized as follows.
@@
  • The CDIS is designed to learn the structural and spatial invariance by integrating channel-wise local and global structural information, and performing joint interaction learning along the different dimensions of the features. It is utilized to construct a shared feature space with better consistency and compactness for subsequent unsupervised learning.
@@ II-A Supervised Cross-view Geo-localization
Cross-view geo-localization (CVGL) is widely recognized as an image retrieval problem based on heterogeneous visual data. However, challenges arise from geometric-semantic heterogeneity and feature distribution shifts across different viewpoints. In traditional CVGL sub-tasks, such as ground-view geo-localization [Liu_2019_CVPR, Hu, Shi_Yu_Liu_Zhang_Li_2020], many researchers have employed explicit geometric alignment methods [shi2019spatial, zhang2023cross, 10601183] to enforce structural consistency at varying viewpoints, while others [regmi2019bridging, Toker_2021_CVPR] have used GANs [goodfellow2014generative] to map source domain images to the target domain. Moreover, in drone-view geo-localization, which is another core sub-task in CVGL, PCL [tian2021uav] attempted to align aerial and satellite images using perspective projection transformation. It also employed GAN to synthesize satellite-style images in order to reduce spatial discrepancies. But it was difficult to handle non-rigid deformations caused by dynamic UAV viewpoints and remained constrained by the synthetic-to-real domain gap. Recently, researchers [ge2024multibranch, lv2024direction] have begun to model latent geometric structures or spatial consistency to enhance the robustness of DVGL. Nevertheless, these methods often focused solely on optimizing geometric structures or spatial representations while neglecting their synergistic relationships. This results in limited generalization under scale variations and spatial shifts in the DVGL. Furthermore, these methods still rely on large amounts of paired data for training, which restricts their applicability in open environments.
@@
Inspired by the aforementioned supervised learning methods, our method utilizes a small amount of paired data for initialization , and further exploits structural and spatial consistency under limited supervision to learn robust and generalizable cross-view representations. Compared with these supervised methods, our method significantly reduces the reliance on paired data while achieving superior performance.
II-B Unsupervised Cross-Domain Retrieval
@@
In recent years, cross-domain image retrieval based on the supervised methods has made significant progress [9963608, 10897882]. CSCL [10812860] enhances cross-domain pedestrian matching by jointly leveraging structural alignment and semantic alignment to extract modality-invariant semantic cues. CSDN [10855601] mitigates the modality discrepancies by integrating visual features with high-level semantic information captured by CLIP [radford2021learning] to enrich the cross-domain representation space. To address the severe performance degradation caused by occlusion, OAFR [10314802] simulates diverse occlusion patterns to enable the model to perceive OTP and PTP occlusions, and further reconstructs the occluded query features using the unoccluded gallery features, which leads to more robust cross-domain retrieval performance. However, the reliance on large-scale annotated data has limited its applicability. Fortunately, recent advances in domain adaptation [10962298, liu2025nonconvex, li2022cross] and semi/unsupervised learning [sun2021semi, jiang2022semi, du2021cross] have demonstrated that robust representations can be learned even under weak or missing labels. This research trend provides important insights for extending cross-domain retrieval toward unsupervised settings and has motivated researchers to explore unsupervised cross-domain retrieval methods [10109197, 9783116, 9891821]. Current methods [Chen_2021_ICCV, Xuan_2021_CVPR, dai2022cluster] utilized powerful feature backbone networks for feature initialization, employing clustering algorithms [ester1996density] to generate pseudo-labels and feature queues that provide stable negative samples for optimizing feature learning. However, these methods are primarily designed for homogeneous data with strong correlations, and their effectiveness tends to degrade significantly when applied to heterogeneous scenarios. To address the limitations, some researchers [yang2022augmented, yang2024shallow] have proposed cross-modal person retrieval methods based on dual-path contrastive learning. Nevertheless, these methods still depend on pre-trained person retrieval models for feature extraction during initialization.
@@ III Proposed Method
Problem Formulation. Given a drone-view image set $\mathcal{X}^{d}=\{x_{i}^{d}\}_{i=1}^{N}$ and a satellite-view image set $\mathcal{X}^{s}=\{x_{j}^{s}\}_{j=1}^{M}$, where each drone image $x_{i}^{d}$ is paired with a corresponding satellite image $x_{j}^{s}$ as a positive pair $(x_{i}^{d},x_{j}^{s})$, the objective is to learn a shared feature space $\mathcal{V}$ that minimizes the distance between positive pairs while maximizing the distance between negative pairs. Existing methods heavily rely on a large amount of paired data as positive samples for training to achieve high performance. In contrast, our method fundamentally differs by mitigating this dependency. The following sections provide a detailed explanation of our method.
Figure 3: Pipeline Overview. The proposed network pipeline consists of two sub-networks. The cross-domain invariance sub-network is designed to learn feature representations that exhibit structural invariance and spatial invariance. The cross-domain transfer sub-network utilizes the invariant knowledge learned by the cross-domain invariance sub-network as a feature representation space to further explore the latent correlations of cross-view from unpaired data.
@@ III-A1 Structural Invariant Learning (SIFL) Block
Given a pair of cross-view images $\{x_{i}^{d},x_{j}^{s}\}$, high-dimensional feature representations are first extracted using the ConvNeXt-Base [liu2022convnet] backbone network. The extracted features are represented as $\mathbf{f}_{v}^{u}\in\mathbb{R}^{C\times H\times W}$, where $u\in\{d,s\}$. Specifically, $v$ is defined as follows: when $u=d$, $v=i$ and when $u=s$, $v=j$. To effectively leverage structural features at different viewpoints, the dimensions $H\times W$ are flattened into a single dimension $L$ to obtain the features $\mathbf{f}_{v}^{ul}\in\mathbb{R}^{C\times L}$. $\mathbf{f}_{v}^{ul}$ serves as the input to the SIFL block.
@@
$\mathbf{f}_{v+}^{u}=\mathcal{N}(\varphi^{3}_{q}(\mathbf{f}_{v}^{ul})+\mathbf{f}_{v}^{ul})$ (1)
$\mathbf{f}_{v+}^{u+}=\mathcal{N}(\mathcal{G}(\psi^{1}_{q}(\mathcal{I}(\mathbf{f}_{v+}^{u}))))+\varphi^{3}_{q}(\mathbf{f}_{v}^{ul})+\mathbf{f}_{v}^{ul}.$ (2)
To further enhance local geometric information, we adopt the same feature processing procedure as in Eq. 1 to extract the feature $\mathbf{f}_{v++}^{u+}$. Although $\mathbf{f}_{v++}^{u+}$ contains rich local geometric information, it still lacks global structural features. Therefore, we replace the CWR module described previously with an MLP $\mathcal{P}(\cdot)$ combined with the squeeze-and-excitation (SE) attention [hu2018squeeze]. In this process, the global average pooling $\mathcal{A}(\cdot)$ is first applied within the SE attention to extract the global information at the channel level. The channel information is then compressed through two linear transformation layers $\mathcal{I}^{+}(\cdot)$, followed by the channel attention weights calculated using the sigmoid function $\mathcal{F}$, which are then combined with the global channel information obtained through the MLP. This further enhances the expression of global channel-level information. Finally, a residual connection is used to fuse the rich local channel-level information from the original features with the enhanced global information. This process can be formulated as
$\mathbf{f}_{v++}^{u+}=\mathcal{N}(\varphi^{3}_{q}(\mathbf{f}_{v+}^{u+})+\mathbf{f}_{v+}^{u+})$ (3)
$\mathbf{f}_{v++}^{u++}\!\!=\!\!\mathcal{F}(\mathcal{I}^{+}(\mathcal{A}(\mathbf{f}_{v++}^{u+})))\times\mathcal{P}(\mathbf{f}_{v++}^{u+})+\varphi^{3}_{q}(\mathbf{f}_{v+}^{u+})+\mathbf{f}_{v+}^{u+}.$ (4)

### v4

@@ Abstract
Traditional supervised drone-view geo-localization (DVGL) methods heavily depend on paired training data and encounter difficulties in learning cross-view correlations from unpaired data. Moreover, when deployed in a new domain, these methods require obtaining the new paired data and subsequent retraining for model adaptation, which significantly increases computational overhead. Existing unsupervised methods have enabled to generate pseudo-labels based on cross-view similarity to infer the pairing relationships. However, geographical similarity and spatial continuity often cause visually analogous features at different geographical locations. The feature confusion compromises the reliability of pseudo-label generation, where incorrect pseudo-labels often drive negative optimization. Given these challenges inherent in both supervised and unsupervised DVGL methods, we propose a novel cross-domain invariant knowledge transfer network (CDIKTNet) with limited supervision, whose architecture consists of a cross-domain invariance sub-network (CDIS) and a cross-domain transfer sub-network (CDTS). This architecture facilitates a closed-loop framework for invariance feature learning and knowledge transfer. The CDIS is designed to learn cross-view structural and spatial invariance from a small amount of paired data that serves as prior knowledge. It endows the shared feature space of unpaired data with implicit cross-view correlations at initialization, which alleviates feature confusion. Based on this, CDTS employs dual-path contrastive learning to further optimize each subspace while preserving consistency in a shared feature space. Extensive experiments demonstrate that CDIKTNet achieves state-of-the-art performance under full supervision compared with those supervised methods, and further surpasses existing unsupervised methods in both few-shot initialization and cross-domain initialization settings.
@@ I Introduction
Cross-view geo-localization has emerged as a promising visual localization technique [39]. Early efforts primarily focus on retrieving ground-view images to geo-referenced satellite-view images for ground-view geo-localization [3]. With the widespread adoption of drones in complex scenarios such as emergency response [29] and urban surveys [22], the demand for autonomous localization in GPS-denied environments has been steadily increasing. This has driven more research interests in drone-view geo-localization (DVGL) [9, 45]. However, the dynamic viewpoints of drones result in scale distortions and non-rigid spatial transformations compared with satellite-view images. This pronounced structural and spatial heterogeneity presents challenges for DVGL. In recent years, benefiting from previous research efforts [47, 5], supervised DVGL methods with paired data have made notable progress. These paired data often require manual labeling, which provides an accurate pairing relationship between satellite-view and drone-view images of the same location.
@@
However, as illustrated in Fig. 1 (A(a)) and (B(a)), existing supervised methods [36, 10, 48] strictly rely on a large amount of paired data to learn cross-view correlations. The lack of sufficient paired data leads to a dramatic decline in model performance. When the supervised ratio drops to 0, or in other words, when there is no paired data and all the data are unpaired, the existing supervised methods become ineffective. We mark this situation with a red X in the figures. Furthermore, dynamic drone mission environments often lead to domain shifts, which in turn require repeated efforts to re-establish cross-view correlations before deployment in new domains. It hinders the applicability in scenarios where pairing relationships are difficult to obtain and cross-domain adaptation is required. As shown in Fig. 1 (B(a)), the new domain where all data are unpaired cannot directly employ the trained model from the source domain. Therefore, researchers have increasingly focused on unsupervised methods for DVGL. Recent work [23] constructed a sparse 3D point cloud from multi-view aerial images and rendered pseudo-satellite images with near-orthographic views via 3D Gaussian splatting to realize geometric transformations of drone-view images. These synthetic images are then matched against real satellite-view images to enable unsupervised training. However, this method is sensitive to 3D reconstruction and pose estimation errors, and it is difficult to bridge the inherent cross-view domain gaps. In contrast, another work [24] utilizes a frozen pre-trained backbone and a learnable adapter to align features from different viewpoints in a shared feature space. The shared feature space serves as a unified latent representation in which drone and satellite views can be directly aligned. Cross-view pseudo-labels are generated automatically based on feature similarity as supervised signals to guide training. This method avoids explicit geometric transformations and aims to intrinsically learn cross-view correlations. Nevertheless, as shown in Fig. 2 (a), when visually similar patterns appear at different geographical locations, the resulting feature confusion in the shared feature space often leads to incorrect pseudo-label generation. This ultimately drives the model toward negative optimization.
@@
In this paper, we propose an efficient learning method with limited supervision for DVGL that mitigates the adverse effects of inaccurate pseudo-labels caused by feature confusion, called the cross-domain invariant knowledge transfer network (CDIKTNet). Specifically, the CDIKTNet consists of two core components, including the cross-domain invariance sub-network (CDIS) and the cross-domain transfer sub-network (CDTS). The CDIS utilizes a small amount of paired data to learn latent cross-view structural and spatial invariance, which serves as prior knowledge. As illustrated in Fig. 2(b), it endows the shared feature space of unpaired data with similar implicit cross-view correlations at initialization to alleviate feature confusion. Based on this, the CDTS introduces dual-path contrastive learning to independently optimize subspaces under different views. Despite being optimized separately, these subspaces maintain alignment in the shared feature space due to the shared invariance prior. The main contributions can be summarized as follows.
@@
  • The CDIS is designed to learn the structural and spatial invariance by integrating channel-wise local and global structural information, and performing joint interaction learning along the different dimensions of the features. It is utilized to construct a shared feature space with better compactness for subsequent unsupervised learning.
@@ II-A Supervised Cross-view Geo-localization
Cross-view geo-localization (CVGL) is widely recognized as an image retrieval problem based on heterogeneous visual data. However, challenges arise from geometric-semantic heterogeneity and feature distribution shifts across different viewpoints. In traditional CVGL sub-tasks, such as ground-view geo-localization [27, 18, 38], many researchers have employed explicit geometric alignment methods [37, 55, 1] to enforce geometric consistency at varying viewpoints, while others [35, 43] have used GANs [16] to map source domain images to the target domain, which aims to reduce appearance discrepancies and maintain style consistency. Moreover, in drone-view geo-localization, which is another sub-task in CVGL, PCL [42] attempted to align drone-view and satellite-view images using perspective projection transformation. It also employed GAN to synthesize satellite-style images in order to reduce cross-view discrepancies and enhance both geometric and style consistency. But it was difficult to handle non-rigid deformations caused by dynamic UAV viewpoints and remained constrained by the synthetic-to-real domain gap. Moreover, such explicit feature consistency mappings fail to capture intrinsic cross-view invariance that differs from these environmentally sensitive features [5, 6]. Recently, researchers [15, 32] have begun to model latent structures or spatial invariance to enhance the robustness of DVGL. Nevertheless, these methods often focused solely on optimizing structures or spatial representations while neglecting their synergistic relationships. This results in limited generalization under scale variations and spatial shifts in the DVGL. Furthermore, these methods still rely on large amounts of paired data for training, which restricts their applicability in open environments.
@@
Inspired by the aforementioned supervised learning methods, our method utilizes a small amount of paired data for initialization, and further exploits structural and spatial invariance under limited supervision to learn robust and generalizable cross-view representations. Compared with these supervised methods, our method significantly reduces the reliance on paired data while achieving superior performance.
II-B Unsupervised Cross-view Retrieval
@@
Supervised methods for cross-view retrieval have achieved remarkable progress [13, 20]. Early cross-view retrieval tasks, exemplified by person re-identification [28, 54, 2, 53], along with the current ground-view geo-localization [18, 38] and drone-view geo-localization tasks [5, 15] that have been attracting extensive attention, effectively establish the instance-level feature consistency through aligning directly based on feature similarity in supervised scenarios. However, the reliance on large-scale annotated data has limited its applicability. Fortunately, recent advances in domain adaptation [58, 30, 25] and semi-unsupervised learning [40, 19, 11] have demonstrated that robust representations can be learned even under weak or missing labels. This research trend provides important insights for extending cross-view retrieval toward unsupervised settings and has motivated researchers to explore unsupervised cross-view retrieval methods [46, 41, 34]. Current methods [4, 49, 8] utilized powerful feature backbone networks for feature initialization, employing clustering algorithms [12] to generate pseudo-labels and feature queues that provide stable negative samples for optimizing feature learning. However, these methods are primarily designed for homogeneous data with strong correlations, and their effectiveness tends to degrade significantly when applied to heterogeneous scenarios. To address the limitations, some researchers [52, 50] have proposed cross-modal person retrieval methods based on dual-path contrastive learning. Nevertheless, these methods still depend on pre-trained person retrieval models for feature extraction during initialization.
@@ III Proposed Method
Problem Formulation. Given a drone-view image set $\mathcal{X}^{d}=\{x_{i}^{d}\}_{i=1}^{N}$ and a satellite-view image set $\mathcal{X}^{s}=\{x_{j}^{s}\}_{j=1}^{M}$, where each drone-view image $x_{i}^{d}$ is paired with a corresponding satellite-view image $x_{j}^{s}$ as a positive pair $(x_{i}^{d},x_{j}^{s})$, the objective is to learn a shared feature space $\mathcal{V}$ that minimizes the distance between positive pairs while maximizing the distance between negative pairs. Existing methods heavily rely on a large amount of paired data as positive samples for training to achieve high performance. In contrast, our method fundamentally differs by mitigating this dependency. The following sections provide a detailed explanation of our method.
Figure 3: Pipeline Overview. The proposed network pipeline consists of two sub-networks. The cross-domain invariance sub-network is designed to learn feature representations that exhibit structural invariance and spatial invariance. The cross-domain transfer sub-network utilizes the invariant knowledge learned by the cross-domain invariance sub-network as a feature representation space to further explore the latent cross-view correlations from unpaired data.
@@ III-A1 Structural Invariant Learning (SIFL) Block
Given a pair of cross-view images $\{x_{i}^{d},x_{j}^{s}\}$, high-dimensional feature representations are first extracted using the ConvNeXt-Base [31] backbone network. The extracted features are represented as $\mathbf{f}_{v}^{u}\in\mathbb{R}^{C\times H\times W}$, where $u\in\{d,s\}$. Specifically, $v$ is defined as follows: when $u$ is $d$ , $v$ is ii and when $u$ is $s$, $v$ is $j$. To effectively leverage structural features at different viewpoints, the dimensions $H\times W$ are flattened into a single dimension $L$ to obtain the features $\mathbf{f}_{v}^{ul}\in\mathbb{R}^{C\times L}$. $\mathbf{f}_{v}^{ul}$ serves as the input to the SIFL block.
@@
$\mathbf{f}_{v+}^{u}=\mathcal{N}\left(\varphi^{3}_{q}(\mathbf{f}_{v}^{ul})+\mathbf{f}_{v}^{ul}\right)$ (1)
$\mathbf{f}_{v+}^{u+}=\mathcal{N}\left(\mathcal{G}\left(\psi^{1}_{q}\left(\mathcal{I}\left(\mathbf{f}_{v+}^{u}\right)\right)\right)\right)+\varphi^{3}_{q}\left(\mathbf{f}_{v}^{ul}\right)+\mathbf{f}_{v}^{ul}$ (2)
To further enhance local geometric information, we adopt the same feature processing procedure as in Eq. 1 to extract the feature $\mathbf{f}_{v++}^{u+}$. Although $\mathbf{f}_{v++}^{u+}$ contains rich local geometric information, it still lacks global structural features. Therefore, we replace the CWR module described previously with an MLP $\mathcal{P}(\cdot)$ combined with the squeeze-and-excitation (SE) attention [17]. In this process, the global average pooling $\mathcal{A}(\cdot)$ is first applied within the SE attention to extract the global information at the channel level. The channel information is then compressed through two linear transformation layers $\mathcal{I}^{+}(\cdot)$, followed by the channel attention weights calculated using the sigmoid function $\mathcal{F}$, which are then combined with the global channel information obtained through the MLP. This further enhances the expression of global channel-level information. Finally, a residual connection is used to fuse the rich local channel-level information from the original features with the enhanced global information. This process can be formulated as
$\mathbf{f}_{v++}^{u+}=\mathcal{N}\left(\varphi^{3}_{q}\left(\mathbf{f}_{v+}^{u+}\right)+\mathbf{f}_{v+}^{u+}\right)$ (3)
$\mathbf{f}_{v++}^{u++}\!\!=\!\!\mathcal{F}\left(\mathcal{I}^{+}\left(\mathcal{A}\left(\mathbf{f}_{v++}^{u+}\right)\right)\right)\times\mathcal{P}\left(\mathbf{f}_{v++}^{u+}\right)+\varphi^{3}_{q}\left(\mathbf{f}_{v+}^{u+}\right)+\mathbf{f}_{v+}^{u+}$ (4)